An object moves 60m due east and then 80 m due north find the distance travelled and the magnitude of displacement


1
Expert's answer
2021-10-24T18:28:58-0400

The distance traveled is the sum of both distance, or


"d=60+80=140\text{ m}."

The magnitude of displacement is


"D=\sqrt{60^2+80^2}=100\text{ m}."